A sample of nitrogen gas occupies a volume of 2.50 L at 25C. If the pressure is held is held constant, what volume does the gas occupy?

1
Expert's answer
2021-09-23T02:37:47-0400

V1/T1 = V2/T2

2.5/298 = V2/368

V2 = 2.5×368/298

V2 = 3.087L